OpenStudy (anonymous):

Together, Steve and John collected 77 baseball cards. John collected 2 more than four times the number Steve collected A. Let x represent the number of cards that Steve has collected. Write an algebraic expression that describes the number of cards that John has collected B. Using the information in Part A and the expression that you wrote, write an algebraic equation to represent the problem C. Solve the equation you wrote in Part B. Show all steps D. How many cards did Steve collect? label answer E. How many cards did John collect? Remember to label answer

OpenStudy (anonymous):

@faylynn

OpenStudy (anonymous):

A. If I understand the given information correctrly then John collected 4 times as much as Steve and 2 more thus 4x + 2. B. x + 4x + 2 = 77 C. 5x + 2 = 77 5x = 75 x = 15 D. Steve collected 15 cards. E. John collected 4 * 15 + 2 = 62 cards. Tiny check 15 + 62 = 77 77 = 77. Solved.

OpenStudy (usman1995):

A. 4x+2 it was easy now i then fancy has given good answer

OpenStudy (anonymous):

@FancyUsername is correct

OpenStudy (anonymous):

yah thats the ans